OpenStudy (anonymous):

can you please help me with x^2+x-20=0

OpenStudy (aroub):

(x+5)(x-4) x=-5 & x=4

OpenStudy (anonymous):

why is it a -5 & a +4?

OpenStudy (aroub):

cos when you factor "x^2+x-20=0" it will become (x+5)(x-4)=0 solve for x , take the first one... x+5=0 , x=-5 and solve for the second one.. x-4=0 , x=4 so x=4 &x=-5 and thats why! :)

OpenStudy (anonymous):

ok, thank you

OpenStudy (aroub):

you're welcome :D

myininaya (myininaya):

\[x^2+5x-4x-20=0\] remember 5x-4x=1x=x so we the polynomial is still the same just written alittle different now factor by grouping first two terms have x in common: x^2+5x=x(x+5) last two terms have -4 in common: -4(x+5) so \[x^2+5x-4x-20=x(x+5)-4(x+5)=0\] x(x+5)-4(x+5)=0 we have two terms now one is x(x+5) and the other is -4(x+5) both of these two terms have the factor (x+5) in common so we can write (x+5)(x-4)=0